OpenStudy (anonymous):

4x-3(5x-8)=23-9(x+2) solve

OpenStudy (anonymous):

-7/2

OpenStudy (anonymous):

sorry -19/2

OpenStudy (anonymous):

19/2

OpenStudy (anonymous):

see here 4x-3(5x-8) = 23-9(x+2) 4x-15x + 24 = 23 - 9x - 18 -11x + 24 = 5 - 9x transpose -11x to right hand side 24 = -9x + 11x 24 = 2x 24/2 = x 12 = x

OpenStudy (anonymous):

albetoimus the answer should be 12

OpenStudy (anonymous):

i got -19/2 where did i go wrong

OpenStudy (anonymous):

-3 * -8 would become 24 not -24

OpenStudy (anonymous):

-11x + 24 = 5 - 9x

OpenStudy (anonymous):

Kindly solve this .you did a mistake

OpenStudy (anonymous):

yes one thing i did wrong here : 24 = -9x + 11x + 5 19 = 2x 19/2 = x

OpenStudy (anonymous):

-2x= -19 => x = 19/2

OpenStudy (anonymous):

4x-3(5x-8)=23-9(x+2) 4x-15x+24=23-9x-18 -11x+24=-9x+5 -2x=-19 x=9.5

OpenStudy (anonymous):

Nice team work.

OpenStudy (anonymous):

ok got it